OpenStudy (anonymous):

subtract 2-y/y-4 - 7y-3/4-y subtract 2-y/y-4 - 7y-3/4-y @Mathematics

OpenStudy (turingtest):

\[{ 2-y \over y-4} - {7y-3 \over 4-y }\]is the problem?

OpenStudy (anonymous):

yes is the answer 2/3?

OpenStudy (turingtest):

well this is not equal to anything, so it can only be simplified:\[{ 2-y \over y-4} - {7y-3 \over 4-y }={2-y \over y-4}+{7y-3 \over y-4}={6y-1 \over y+4}\]

OpenStudy (turingtest):

sorry should be y-4 on the bottom of the last step

OpenStudy (turingtest):

\[{6y-1 \over y-4}\]

OpenStudy (anonymous):

so i just add the two together? how did you figure out the y-4?

OpenStudy (turingtest):

all I did was take the negative of the second denominator to get both the same -(4-y)=y-4 Note this also turned the "-" sign into a "+" in the middle

OpenStudy (turingtest):

i.e we basically multiplied the second expression by (-1/-1)

OpenStudy (anonymous):

so would a shorter way be just to multiply by -1/-1

OpenStudy (turingtest):

same thing

OpenStudy (anonymous):

ok thanks. im checking my work before i turn it in.